OpenStudy (anonymous):

will give medals but pleaseeee help. Solve for x: 4 − (x + 2) < −3(x + 4) x < −7 x > −7 x < −9 x > −9

OpenStudy (anonymous):

the first one

OpenStudy (calculusxy):

This is basically solving for x in a regular equation. So what you would just do is instead of having the less than symbol, you would place the equal sign.

OpenStudy (anonymous):

@calculusxy wait ill do the work and show you where i am stuck at

OpenStudy (calculusxy):

Okay

OpenStudy (anonymous):

\[ 4-(x+2)<-3x-6\]

OpenStudy (anonymous):

@calculusxy ^^^^

OpenStudy (anonymous):

i don't know how to do with the other side do i take away the brackets and flip the sign of 4 to + and add the like terms ?

OpenStudy (anonymous):

after the negative sign there is a 1 so multiply negative one by x and 2

OpenStudy (anonymous):

so it would be -1x -2 ? @joquez14

OpenStudy (calculusxy):

Just like what @joquez14 said. what we can think is that before the parentheses, there is a -1 (negative because of the minus sign). \[4-1(x+2) <-3x + 12\] that means that you would have to do \[-1 \times x = -x\]and \[-1 \times 2 = -2\]

OpenStudy (calculusxy):

Now we can simplify it even further wit those two terms. \[-4 -x-2<-3x-12\]

OpenStudy (calculusxy):

Sorry i meant -12 not +12

OpenStudy (anonymous):

okay so we add together -4 and -2 @calculusxy

OpenStudy (calculusxy):

yes to simplify that side

OpenStudy (calculusxy):

combine the like terms to make solving this easier

OpenStudy (anonymous):

okay so we add the x's

OpenStudy (calculusxy):

so this is where you need to really cancel out an x from a side.

OpenStudy (anonymous):

so it would be 6<-2x

OpenStudy (calculusxy):

\[-4-x-2=-3x - 12\] \[-6 - x=-3x - 12\] \[-6 =-2x - 12\]

OpenStudy (calculusxy):

yes you're correct

OpenStudy (anonymous):

but then the answer would 4 and none of my answer choices are 4 ? did i do an error ?

OpenStudy (anonymous):

****-3 i mean

OpenStudy (calculusxy):

what are your answer choices then?

OpenStudy (anonymous):

they are at the top

OpenStudy (calculusxy):

oh sorry :\

OpenStudy (anonymous):

its not supposed to ve - 2x i thought i think its supposed to be - 4x

OpenStudy (calculusxy):

@joquez14 if we did -3 + 1 it would equal to -2

OpenStudy (anonymous):

yup thats what I'm getting -2

OpenStudy (anonymous):

why you add one x was negative so u souldnt have added an positive 1

OpenStudy (anonymous):

-1x+-3x ? @joquez14

OpenStudy (anonymous):

ohhh -4

OpenStudy (anonymous):

my point exactly

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

aren't you suppose to do the inverse

OpenStudy (anonymous):

im not to sure since were combining like terms

OpenStudy (anonymous):

I wish I could give you all medals. You all have the right idea, it just looks like you got mixed up with the combining 4 − (x + 2) < −3(x + 4) Start by distributing 4 - x - 2 < -3x - 12 Combine like terms on the left 2 - x < -3x - 12 Add 3x to both sides 2 + 2x < -12 Subtract 2 from both sides 2x < -14 Divide both sides by 2 x < -7

OpenStudy (anonymous):

oh yeah ur right

OpenStudy (anonymous):

thank you so much guys literally all of you have been such a tremendous helppppppppp @joquez14 @calculusxy and thanks for clearing it up @peachpi

OpenStudy (anonymous):

np